### v2

# Hardness of Regular Expression Matching with Extensions

Hardness of Regular Expression Matching with Extensions
Taisei Nogami
Waseda University
sora410@fuji.waseda.jp
   Tachio Terauchi
Waseda University
terauchi@waseda.jp
The regular expression matching problem asks whether a given regular expression of length $m$ matches a given string of length $n$. As is well known, the problem can be solved in $O(nm)$ time using Thompson’s algorithm. Moreover, recent studies have shown that the matching problem for regular expressions extended with a practical extension called lookaround can be solved in the same time complexity. In this work, we consider four well-known extensions to regular expressions called backreference, squaring, intersection and complement, and we show that, unlike in the case of lookaround, the matching problem for regular expressions extended with any of the four (for backreference, even when restricted to one capturing group) cannot be solved in $O(n^{2-\varepsilon}\operatorname{poly}(m))$ time for any constant $\varepsilon>0$ under the Orthogonal Vectors Conjecture. Moreover, we study the matching problem for regular expressions extended with complement in more detail, which is also known as extended regular expression (ERE) matching. We show that there is no ERE matching algorithm that runs in $O(n^{\omega-\varepsilon}\operatorname{poly}(m))$ time ($2\leq\omega<2.3714$ is the exponent of square matrix multiplication) for any constant $\varepsilon>0$ under the $k$-Clique Hypothesis, and there is no combinatorial ERE matching algorithm that runs in $O(n^{3-\varepsilon}\operatorname{poly}(m))$ time for any constant $\varepsilon>0$ under the Combinatorial $k$-Clique Hypothesis. This shows that the $O(n^{3}m)$-time algorithm introduced by Hopcroft and Ullman in 1979 and recently improved by Bille et al. to run in $O(n^{\omega}m)$ time using fast matrix multiplication was already optimal in a sense, and sheds light on why the theoretical computer science community has struggled to improve the time complexity of ERE matching with respect to $n$ and $m$ for more than 45 years.
1 Introduction
Regular expressions, originally introduced by Kleene [37], are one of the most successful concepts in computer science, both in theory and practice. Among problems related to regular expressions, the regular expression matching problem is of particular importance, which asks whether a given regular expression $r$ of length $m$ matches a given string $w$ of length $n$. As is well known, the problem can be solved in $O(nm)$ time using an algorithm introduced by Thompson [55]. Roughly, it converts a given regular expression $r$ to an equivalent nondeterministic finite automaton (NFA) and simulates it on a given string $w$.
On the other hand, regular expression matching becomes much harder when extensions are present. A prominent example is backreference, an extension that allows reusing previously matched substrings. The extension is practically popular and supported in the standard libraries of many popular programming languages including Java, Python, JavaScript and more. Unfortunately, we cannot solve the matching problem for regular expressions extended with backreference (rewbs for short) using Thompson’s algorithm because rewbs are strictly more expressive than regular languages (in fact, they are not even context-free) [11] and therefore an equivalent NFA does not in general exist. Instead, it can be solved by simulating memory automata, which are a model equivalent to rewbs introduced by Schmid [51]. Nevertheless, the resulting matching algorithm for rewbs with at most $\ell$ capturing groups takes $O(n^{2\ell+1}m)$ time [51, 52, 45], and at present, even when $\ell=1$, we incur an additional $n^{2}$ factor compared to the running time of Thompson’s algorithm. Another practically popular extension is bounded repetition $r\{i,j\}$, which denotes repeated concatenation of expression $r$ by specifying a lower bound $i$ and an upper bound $j$ on the number of repetitions. In the special case $i=j=2$, the extension is called squaring [40]. Although bounded repetition can be eliminated by unrolling the repetition bounds, the resulting expression can become exponentially larger. Therefore, the naive algorithm eliminating it and running Thompson’s algorithm takes at least exponential time in $m$.
Other well-known extensions that make the matching problem difficult are intersection and complement, where regular expressions extended with each of them are called semi-extended regular expressions (semi-EREs) and extended regular expressions (EREs), respectively. We note that EREs essentially subsume semi-EREs because intersection can be expressed using complement and alternation with only a linear increase in the length of the expression. The best known time complexity of both semi-ERE matching and ERE matching with respect to $n$ and $m$ is $O(n^{\omega}m)$ [15], where $2\leq\omega<2.3714$ [7] is the exponent of square matrix multiplication.
In light of the above, we pose the following question: Do these extensions actually increase the time complexity of regular expression matching, or do they only appear hard because we have not yet discovered the right algorithmic ideas? Notably, lookaround is a practically popular extension to regular expressions like backreference and bounded repetition. The extension can be used to test surrounding contexts of a position and is often used to mimic intersection and complement. Recent studies have shown a surprising fact that this extension does not actually increase the complexity of regular expression matching by presenting $O(nm)$-time algorithms for the matching problem with the extension [39, 24, 10].
1.1 Result I: Hardness of Matching with Extensions under OVC
In this paper, we consider the four extensions described above, backreference, squaring, intersection and complement, and show that regular expression matching extended with each of the four (for backreference, even when restricted to one capturing group) is hard under the Orthogonal Vectors Conjecture (OVC). We note that this also holds under the Strong Exponential Time Hypothesis (SETH) [33] because SETH implies OVC [56, 58].
Theorem 1.
Unless OVC fails, there is no algorithm that solves rewb matching in $O(n^{2-\varepsilon}\operatorname{poly}(m))$ time for any constant $\varepsilon>0$, even when restricted to one capturing group. [^1] Here, $\operatorname{poly}(m)$ denotes a constant-degree polynomial in $m$.
We refer to regular expressions extended with squaring as RSQs.
Theorem 2.
Unless OVC fails, there is no algorithm that solves RSQ matching in $O(n^{2-\varepsilon}\operatorname{poly}(m))$ time for any constant $\varepsilon>0$.
Theorem 3.
Unless OVC fails, there is no algorithm that solves ERE matching in $O(n^{2-\varepsilon}\operatorname{poly}(m))$ time for any constant $\varepsilon>0$, even when restricted to semi-EREs.
Theorems 1, 2 and 3 state a much stronger statement than merely that we should not hope for a matching algorithm for rewbs, RSQs or semi-EREs that runs in $O(nm)$ time. Indeed, it states that if we want to solve rewb matching, RSQ matching or semi-ERE matching in truly subquadratic time in $n$ (i.e., $n^{2-\varepsilon}$ for some $\varepsilon>0$), then we are forced to admit a superpolynomial blowup in $m$ unless OVC fails.
The results are of interest not only to theorists but also to practitioners. They provide evidence that the real-world regular expression engines which guarantee worst-case $O(nm)$-time complexity (e.g., that in the de facto standard regular expression library for Rust [^2] The engine guarantees worst-case $O(nm)$-time complexity, although it supports bounded repetition (and hence also squaring) [1]. This is consistent with Theorem 2 (without refuting OVC) because here $m$ denotes the length of the input regular expression after bounded repetition has been eliminated. See the “Untrusted input” section in [1].) are all forced to give up the full support for backreference, squaring and intersection, and would need to somehow restrict their usage such that our reductions in the proofs of Theorems 1, 2 and 3 cannot apply. Thus, the results are of great practical value because they provide guidelines on how far the expressiveness can be increased in regular expression engines while retaining the good complexity.
We also observe that Theorem 3 shows the hardness of the matching problem extended with intersection relative to that extended with lookaround, which is solvable in $O(nm)$ time as mentioned above. It is known that lookaround is related to intersection and can express certain restricted forms of it (see, e.g., [39]). Theorem 3 states that, nevertheless, intersection makes the problem strictly more difficult than lookaround does, or OVC fails.
1.2 Result II: Hardness of Matching with Complement under $k$-Clique Hypotheses
We then study the ERE matching problem in more detail. The line of research on improving its time complexity dates back to 1979, when Hopcroft and Ullman introduced an ERE matching algorithm that runs in $O(n^{3}m)$ time [29]. Until 2025, apart from the word-size speedup due to Yamamoto and Miyazaki [62], this had remained the best known time complexity with respect to $n$ and $m$ for both semi-ERE matching and ERE matching. In 2025, Bille et al. made significant progress by showing that Hopcroft and Ullman’s algorithm can run in $O(n^{\omega}m)$ time using fast matrix multiplication [15].
We emphasize that, nevertheless, Hopcroft and Ullman’s algorithm remains essentially unsurpassed with respect to $n$ and $m$. In fact, as noted in [15], the improvement from $O(n^{3}m)$ to $O(n^{\omega}m)$ can be obtained by using fast matrix multiplication without substantially changing the original algorithm. Moreover, it is still the best combinatorial algorithm with respect to $n$ and $m$ (except the word-size speedup of [62]). It remained a mystery why the algorithm has stood the test of time for over 45 years.
In this paper, we follow an approach introduced by Abboud et al. [2] to show the hardness of ERE matching under the $k$-Clique Hypothesis and the Combinatorial $k$-Clique Hypothesis. This implies, under the hypotheses, that Hopcroft and Ullman’s algorithm is combinatorially optimal and optimal with fast matrix multiplication in the sense that the exponents of $n$ in the respective running times cannot be shaved by any constant while preserving the running times to be constant-degree polynomial in mm. [^3] We refer to Lemma 14 for why Theorem 4 implies this optimality.
Theorem 4.
Let $\varepsilon>0$ be an arbitrary constant. Unless the $k$-Clique Hypothesis fails, there is no ERE matching algorithm that runs in $O(n^{\omega-\varepsilon}\operatorname{poly}(m))$ time. Moreover, unless the Combinatorial $k$-Clique Hypothesis fails, there is no combinatorial ERE matching algorithm that runs in $O(n^{3-\varepsilon}\operatorname{poly}(m))$ time.
The result sheds some light on the mystery of why Hopcroft and Ullman’s algorithm stood the test of time for such a long time. That is, the algorithm was already optimal in the sense mentioned above except for the use of fast matrix multiplication. We note that if matrix multiplication can be done in quadratic time (i.e., $\omega=2$), Theorem 3 also establishes such optimality of the algorithm using fast matrix multiplication even for semi-ERE matching under OVC.
1.3 Technical Overview
All lower bounds shown in this paper are of the form $O(n^{c-\varepsilon}\operatorname{poly}(m))$ and therefore do not rule out algorithms that run in superpolynomial time in $m$ (e.g., $O(n^{c-1}2^{m})$). Still, this polynomial-in-$m$ lower bound rules out algorithms that are efficient with respect to both $n$ and $m$. Moreover, restricting the time complexity of RSQ matching and ERE matching with respect to $m$ is even necessary for deriving a meaningful lower bound on the complexity with respect to $n$ and $m$ because we could improve the complexity to be linear in $n$ if we were to allow the complexity to be unbounded in $m$. Indeed, the naive RSQ matching algorithm mentioned earlier that eliminates squares by unrolling and uses Thompson’s algorithm runs in time linear in $n$ but exponential in $m$. ERE matching can also be solved in linear time in $n$ by converting a given ERE to an equivalent NFA and running Thompson’s algorithm. But this makes the algorithm’s running time nonelementary in $m$ [53]. Rosu has proposed a less extreme ERE matching algorithm that runs in $O(n^{2}\,2^{m})$ time [49].
We prove our theorems by reductions from the respective problems. A key common technique in our reductions is to exploit the polynomial-in-$m$ lower-bound property to allow the regular expressions to be of polylogarithmic lengths with respect to the sizes of the instances to be reduced. For example, in the proofs of Theorems 1, 2 and 3, we show reductions from the Orthogonal Vectors problem, which asks whether given sets $A$ and $B$ containing the same number of bit-vectors of dimension $d$ have a pair of orthogonal vectors $a\in A$ and $b\in B$. We encode the vectors in $A$ and $B$ into a string and construct a regular expression $r$ that nondeterministically guesses two vectors $a\in A$ and $b\in B$ from the string and checks their orthogonality. In proving a lower bound under OVC, we may assume that $d$ is logarithmic in the number of vectors in $A$ and $B$. Therefore, we can let the length of $r$ be polynomial in $d$ for proving our polynomial-in-$m$ lower bounds. As we shall show, such $r$ can be constructed by exploiting this property and the respective regular expression extensions. A similar argument applies to Theorem 4 where we show a reduction from the $3k$-Clique problem.
The rest of the paper is organized as follows. Section 2 discusses related work. Section 3 defines preliminary notions. Sections 4 and 5 contain proofs of the theorems mentioned above. Section 6 concludes the paper with a discussion on future work.
2 Related Work
In this section, to clarify the relation to our work, we express running times from the prior work with respect to the regular expression length $m$ (and the input string length $n$) even when the original bounds were given using more refined parameters of regular expressions, unless stated otherwise. We first discuss the prior work on regular expression matching without extensions. A line of work has shown that the complexity of Thompson’s algorithm can be improved [41, 12, 13, 16, 17, 14]. On the other hand, Backurs and Indyk showed that the problem cannot be solved in $O((nm)^{1-\varepsilon})$ time for any $\varepsilon>0$ unless SETH fails [9]. Later, Abboud and Bringmann introduced a hypothesis, called the Formula-SAT Hypothesis, and showed a more fine-grained lower bound assuming it [3].
Next, we discuss the prior work on rewb matching. Aho showed that the problem is NP-complete [6]. Moreover, it is known that rewbs can be regarded as a generalization of Angluin’s pattern languages (also known as patterns with variables), whose matching problem is already NP-complete [8, 21], even under bounded parameterizations [50, 22]. At this point, one may wonder about the novelty of Theorem 1 because NP-hardness already rules out rewb matching algorithms that run in $O(n^{2-\varepsilon}\operatorname{poly}(m))$ time (in fact, even those that run in time polynomial in $n+m$) assuming $\mathrm{P}\neq\mathrm{NP}$. However, it cannot rule out matching algorithms for rewbs with at most $\ell$ capturing groups for a fixed $\ell$, because the problem is solvable in $O(n^{2\ell+1}m)$ time as stated in the introduction and hence is in $\mathrm{P}$ when $\ell$ is fixed. By contrast, our Theorem 1 can rule out those for rewbs with even one capturing group under OVC. Furthermore, to our knowledge, there is no known implication between OVC and $\mathrm{P}\neq\mathrm{NP}$ (although both follow from SETH [58, 20]) and therefore these hardness results appear to be independent. Regarding upper bounds for the problem, Freydenberger and Schmid introduced a deterministic variant of rewbs together with algorithms that decide whether a given rewb is deterministic in their sense in $O(m^{2})$ time and solve the matching problem for those rewbs in $O(m^{2}+n\ell)$ time [23]. Therefore, combining this with our Theorem 1, we obtain evidence under OVC that their determinism makes rewb matching strictly more tractable (even when $\ell=1$). Terauchi [54] has shown that the matching problem for rewbs satisfying a condition called constant-degree ambiguity (CDA) can be solved in linear time in $n$. They leave open the decidability of checking whether a given rewb satisfies CDA. Additionally, Nogami and Terauchi [45] recently proposed an $O(n^{2}m^{2})$-time algorithm for a certain subclass of rewbs with one capturing group, improving the previously known best $O(n^{3}m)$ algorithm for the subclass with respect to $n$. We comment on the expressive power of rewbs. As mentioned in the introduction, backreference differs from bounded repetition, intersection and complement in that it strictly increases the expressive power of regular expressions. The expressive power of rewbs is known to be incomparable with that of context-free languages (CFLs) [18, 11]. Furthermore, recent papers have shown that it is properly contained in that of EDT0L languages and incomparable with both the class of stack languages and that of multiple context-free languages (MCFLs), while a syntactic condition of rewbs, called the closed-star condition, decreases its expressive power to be properly contained in subclasses of those, namely the class of nonerasing stack languages and that of unary-MCFLs [43, 44, 46].
Next, we discuss the prior work on the matching problem extended with bounded repetition. Kilpeläinen and Tuhkanen showed that the problem is in P [35]. The problem for subclasses of bounded repetition has been extensively studied. Indeed, various formalizations of determinism have been introduced in the literature [36, 30, 34, 31, 25, 19, 27]. One such formalization is strong determinism. Whether a given regular expression with bounded repetition is strongly deterministic can be decided in linear time [19, 27], and the matching problem for strongly-deterministic expressions can be solved in $O(n\log{\log{m}})$ time after a preprocessing phase that runs in $O(m)$ expected time [27]. Glaunec et al. [26] proposed a matching algorithm for a certain subclass with no nested bounded repetition that runs in $O(nm)$ time and with the running time independent of the repetition bounds. Therefore, as in the case of deterministic rewb matching, these prior results and our Theorem 2 provide evidence under OVC that the matching problems for the respective subclasses considered by the prior works are strictly more tractable. [^4] For strongly-deterministic expressions, we strengthen OVC so that it rules out even randomized algorithms. Such a variant can be found in the literature [59]. Also, [27] notes that their preprocessing can be made deterministic. For RSQs, Meyer and Stockmeyer showed that deciding universality (and hence also equivalence) requires exponential space and therefore it is not in P [40]. It is also known that RSQ matching extended with intersection is in LOGCFL [47] and that extended with complement is in P [53].
Next, we discuss the prior work on ERE matching. As mentioned in the introduction, Hopcroft and Ullman [29] proposed in 1979 an algorithm that runs in $O(n^{3}m)$ time and no polynomial improvement over this with respect to $n$ and $m$ had existed for many years prior to Bille et al. [15]. ERE matching is extensively studied and has an interesting history. In fact, many papers had claimed to propose an improved $O(n^{2}m)$-time algorithm [28, 60, 38, 32] ([60] considers semi-ERE matching). However, all of these claims were later found to be incorrect; see [48, 49] for more details. Additionally, a number of studies had improved the complexity of ERE matching with respect to more refined parameters than just $n$ and $m$ [61, 62, 15]. As also mentioned in the introduction, Bille et al. [15] recently adopted fast matrix multiplication to improve the time and space complexity of Hopcroft and Ullman’s algorithm and that of the algorithm by Yamamoto and Miyazaki [62]. Furthermore, as mentioned in Section 1.3, Rosu proposed an algorithm that runs in time quadratic in $n$, but exponential in $m$ [49]. On the hardness side, Petersen proved that semi-ERE matching is LOGCFL-complete by showing a reduction $f$ from every context-free language (CFL) to semi-ERE matching [47]. Combining $f$, which is combinatorial, and the combinatorial reduction from $3k$-Clique to CFL recognition shown by Abboud et al. [2], we can derive the following hardness of semi-ERE matching: (i) there is no semi-ERE matching algorithm that runs in $O((n+m)^{\omega-\varepsilon})$ time for any $\varepsilon>0$ under the $k$-Clique Hypothesis, and (ii) there is no combinatorial semi-ERE matching algorithm that runs in $O((n+m)^{3-\varepsilon})$ time for any $\varepsilon>0$ under the Combinatorial $k$-Clique Hypothesis. However, note that these bounds cannot establish the optimality of Hopcroft and Ullman’s algorithm in the sense described in Section 1.2. By contrast, our Theorem 4 can do it. This is because our reduction outputs EREs of length only polylogarithmic in the input size, whereas $f$ outputs EREs of length linear in the input size.
Finally, we note that our form of multivariate lower bounds is not new. For example, Abboud et al. [5] proved that under OVC, for any $\varepsilon>0$, there is no $n^{2-\varepsilon}\,2^{o(k)}$-time algorithm that decides if the diameter of a given graph of treewidth $k$ is no more than $2$.
3 Preliminaries
For a string or a vector $s$, we write $s^{R}$ for the reverse of $s$. We usually omit the string concatenation symbol, but occasionally write it explicitly using $\odot$. The syntax of regular expressions over an alphabet $\Sigma$ is given by the following grammar:
$r::=a\mid\lambda\mid\emptyset\mid r_{1}r_{2}\mid r_{1}+r_{2}\mid r^{*}$
where $a\in\Sigma$ and $\lambda$ denotes the empty string. For brevity, we use the character set notation $a_{1}a_{2}\cdots a_{k}]$, which is a shorthand for the regular expression $a_{1}+a_{2}+\cdots+a_{k}$ where $a_{1},\dots,a_{k}\in\Sigma$. In particular, we regard $\Sigma$ as the character set consisting of all elements of $\Sigma$. As standard, for a string or a regular expression $\alpha$ and a non-negative integer $i$, we write $\alpha^{i}$ for $\alpha$ concatenated $i$ times. The language of a regular expression is defined in the standard way. A regular expression $r$ matches a string $w$ if $w$ is an element of the language of $r$, which is also written as $w\in r$ by a slight abuse of notation. Two regular expressions are equivalent if they define the same language.
4 OV to Semi-ERE, Rewb, and RSQ Matching
In this section, we prove Theorems 1, 2 and 3. The Orthogonal Vectors (OV) problem is the following: given sets $A,B\subseteq\{0,1\}^{d}$ containing the same number $n$ of bit-vectors of dimension $d$, decide if there are two vectors $a\in A$ and $b\in B$ such that $a$ and $b$ are orthogonal. A naive algorithm solves OV in $O(n^{2}d)$ time.
Conjecture (Orthogonal Vectors Conjecture, see [57]).
For every constant $\varepsilon>0$, there is a constant $c\geq 1$ such that OV cannot be solved in $O(n^{2-\varepsilon})$ time on instances with $d=c\log{n}$.
As mentioned in the introduction, it is known that SETH implies OVC (see Lemma A.1 in [58] for the proof), and therefore any hardness result under OVC also holds under SETH. Moreover, OVC is supported by evidence beyond its implication from SETH. See, for example, [57] for more details.
4.1 Semi-ERE matching
For simplicity, we prove Theorem 3 before Theorems 1 and 2. Semi-ERE [29] extends regular expressions with the intersection construct $r_{1}\cap r_{2}$. The language of a semi-ERE is defined by letting the language of $r_{1}\cap r_{2}$ be the intersection of the languages of $r_{1}$ and $r_{2}$. For example, $(aa)^{*}\cap(aaa)^{*}$ is equivalent to $(aaaaaa)^{*}$, whose language is $\{a^{6n}\mid n\geq 0\}$. The following lemma implies Theorem 3.
Lemma 5.
If semi-ERE matching on a string of length $n$ and a semi-ERE of length $m$ can be solved in $O(n^{2-\varepsilon}\operatorname{poly}(m))$ time for some $\varepsilon>0$, then OV on sets of $n$ vectors of dimension $d$ can be solved in $O(n^{2-\varepsilon^{\prime}}\operatorname{poly}(d))$ time for some $\varepsilon^{\prime}>0$ and OVC is false.
Proof.
We construct a reduction from OV to semi-ERE matching as follows. Given sets of vectors $A=\{a_{1},\dots,a_{n}\}$ and $B=\{b_{1},\dots,b_{n}\}$, we encode them into a string $w$ in the following way:
$w=\textup{\#}a_{1}\textup{\#}a_{2}\cdots\textup{\#}a_{n}\textup{\textdollaroldstyle}b_{1}^{R}\textup{\#}b_{2}^{R}\textup{\#}\cdots b_{n}^{R}\textup{\#}.$
Both the length of $w$ and the time required for this construction are $O(nd)$. We then define a semi-ERE $r$ over the alphabet $\Sigma=\{0,1,\textup{\#},\textup{\textdollaroldstyle}\}$ as follows:
$r$ $=[01\textup{\#}]^{*}\textup{\#}(r_{\mathrm{zero}}\cap[01]r_{\mathrm{zero}}[01]\cap\cdots\cap[01]^{d-1}r_{\mathrm{zero}}[01]^{d-1})\textup{\#}[01\textup{\#}]^{*},$
$r_{\mathrm{zero}}$ $=0\Sigma^{*}[01]+1\Sigma^{*}0.$
Both the length of $r$ and the time required for this construction are $O(d^{2})$.
We claim that there exist vectors $a_{i}\in A$ and $b_{j}\in B$ that are orthogonal if and only if $r$ matches $w$. For the only if direction, we can decompose $w$ as $w_{1}\textup{\#}a_{i}w_{2}b_{j}^{R}\textup{\#}w_{3}$ where $a_{i}\in A$ and $b_{j}\in B$ are orthogonal, $w_{1},w_{3}\in[01\textup{\#}]^{*}$ and $w_{2}\in\Sigma^{*}$. Thus, it remains to show that every subexpression $01]^{k}r_{\mathrm{zero}}[01]^{k}$ of $r$ matches $a_{i}w_{2}b_{j}^{R}$ for $k\in\{0,\dots,d-1\}$, and this holds because $a_{i}$ and $b_{j}$ are orthogonal. Conversely, suppose that $r$ matches $w$. The delimiter \$ enforces that, there exist $a_{i}\in A,b_{j}\in B$ and $w_{1},w_{2}\in[01\textup{\#}]^{*}$ such that every subexpression $01]^{k}r_{\mathrm{zero}}[01]^{k}$ $(k\in\{0,\dots,d-1\})$ matches the substring $a_{i}w_{1}\textup{\textdollaroldstyle}w_{2}b_{j}^{R}$. By the definition of $r_{\mathrm{zero}}$, it follows that $a_{i}$ and $b_{j}$ are orthogonal.
Without loss of generality, we may assume that $2-\varepsilon\geq 1$. Using the semi-ERE matching algorithm in the hypothesis, we can solve the OV problem with $d=c\log{n}$ for any constant $c\geq 1$ in time
$O(nd+d^{2}+(nd)^{2-\varepsilon}\operatorname{poly}(d^{2}))=O(n^{2-\varepsilon}\operatorname{poly}(d))=O(n^{2-\varepsilon/2}).\qed$
Remark 6.
[…]
4.2 Rewb matching
[…]
Lemma 7.
If 1-rewb matching on a string of length $n$ and a 1-rewb of length $m$ can be solved in $O(n^{2-\varepsilon}\operatorname{poly}(m))$ time for some $\varepsilon>0$, then OV on sets of $n$ vectors of dimension $d$ can be solved in $O(n^{2-\varepsilon^{\prime}}\operatorname{poly}(d))$ time for some $\varepsilon^{\prime}>0$ and OVC is false.
Proof.
We construct a reduction from OV to 1-rewb matching as follows. Given sets of vectors $A=\{a_{1},\dots,a_{n}\}$ and $B=\{b_{1},\dots,b_{n}\}$, we encode them into a string $w$ in the following way:
$w=(s\textup{\textcentoldstyle})^{d+1}\text{ where }s=a_{1}\textup{\#}a_{2}\textup{\#}\cdots a_{n}\textup{\#}\textup{\textdollaroldstyle}\textup{\#}b_{1}^{R}\textup{\#}b_{2}^{R}\cdots\textup{\#}b_{n}^{R}.$
[…]
$r$ $=[01\textup{\#}]^{*}(\textup{\#}\Sigma^{*}\textup{\#})_{1}[01\textup{\#}]^{*}\textup{\textcentoldstyle}r_{\mathrm{zero},1}\textup{\textcentoldstyle}r_{\mathrm{zero},2}\textup{\textcentoldstyle}\cdots\textup{\textcentoldstyle}r_{\mathrm{zero},d}\textup{\textcentoldstyle},$
$r_{\mathrm{zero},k}$ $=[01\textup{\#}]^{*}(0[01]^{k-1}\backslash 1[01]^{k-1}[01]+1[01]^{k-1}\backslash 1[01]^{k-1}0)[01\textup{\#}]^{*}\text{ for }k\in\{1,\dots,d\}.$
Both the length of $r$ and the time required for this construction are $O(d^{2})$.
We claim that there exist vectors $a_{i}\in A$ and $b_{j}\in B$ that are orthogonal if and only if $r$ matches $w$. For the only if direction, we can decompose $w$ as
$w=w_{0}a_{i}\textup{\#}w_{1}\textup{\textdollaroldstyle}w_{2}\textup{\#}b_{j}^{R}w_{3}\textup{\textcentoldstyle}(w_{0}a_{i}\textup{\#}w_{1}\textup{\textdollaroldstyle}w_{2}\textup{\#}b_{j}^{R}w_{3}\textup{\textcentoldstyle})^{d}$
where $a_{i}\in A$ and $b_{j}\in B$ are orthogonal, and $w_{0},w_{1},w_{2},w_{3}\in[01\textup{\#}]^{*}$. We let $r$ capture $\textup{\#}w_{1}\textup{\textdollaroldstyle}w_{2}\textup{\#}$ in its capturing group. By splitting $w$ using the delimiter ¢, it suffices to show that for each $k\in\{1,\dots,d\}$, either (i) $a_{i}\in[01\textup{\#}]^{*}0[01]^{k-1}$ and $b_{j}^{R}\in[01]^{k-1}[01][01\textup{\#}]^{*}$ or (ii) $a_{i}\in[01\textup{\#}]^{*}1[01]^{k-1}$ and $b_{j}^{R}\in[01]^{k-1}0[01\textup{\#}]^{*}$ holds. This indeed holds because $a_{i}$ and $b_{j}$ are orthogonal.
Conversely, suppose that $r$ matches $w$. The delimiters ¢ and \$ enforce that the subexpression $01\textup{\#}]^{*}(\textup{\#}\Sigma^{*}\textup{\#})_{1}[01\textup{\#}]^{*}$ matches $s$ in such a way that the capturing group matches $\textup{\#}w_{1}\textup{\textdollaroldstyle}w_{2}\textup{\#}$ for some $w_{1},w_{2}\in[01\textup{\#}]^{*}$. By the construction of $w$, there are vectors $a_{i}\in A$ and $b_{j}\in B$ such that $s=w_{0}a_{i}\textup{\#}w_{1}\textup{\textdollaroldstyle}w_{2}\textup{\#}b_{j}^{R}w_{3}$ for some $w_{0},w_{3}\in[01\textup{\#}]^{*}$. Then, every subsequent subexpression $r_{\mathrm{zero},1},\dots,r_{\mathrm{zero},d}$ matches $s$ so that $\backslash 1$ matches $\textup{\#}w_{1}\textup{\textdollaroldstyle}w_{2}\textup{\#}$. This implies that $a_{i}$ and $b_{j}$ are orthogonal.
Without loss of generality, we may assume that $2-\varepsilon\geq 1$. Using the 1-rewb matching algorithm in the hypothesis, we can solve the OV problem with $d=c\log{n}$ for any constant $c\geq 1$ in time
$O(nd^{2}+d^{2}+(nd^{2})^{2-\varepsilon}\operatorname{poly}(d^{2}))=O(n^{2-\varepsilon}\operatorname{poly}(d))=O(n^{2-\varepsilon/2}).\qed$
4.3 RSQ matching
In what follows, we prove Theorem 2 by establishing the following Lemma 8. RSQ [40] extends regular expressions with the squaring construct ${r}\{2\}$, which denotes an expression equivalent to $rr$.
Lemma 8.
If RSQ matching on a string of length $n$ and an RSQ of length $m$ can be solved in $O(n^{2-\varepsilon}\operatorname{poly}(m))$ time for some $\varepsilon>0$, then OV on sets of $n$ vectors of dimension $d$ can be solved in $O(n^{2-\varepsilon^{\prime}}\operatorname{poly}(d))$ time for some $\varepsilon^{\prime}>0$ and OVC is false.
Proof.
We construct a reduction from OV to RSQ matching as follows. Let $A$ and $B$ be sets of bit-vectors of dimension $d$ satisfying $|A|=|B|=n$. For $a\in A$, we write $a_{i}$ for the $i$-th coordinate of $a$ and let $\mathrm{VG}_{1}(a)=\textup{\textcentoldstyle}_{1}\overline{a_{1}}\textup{\textcentoldstyle}_{1}\overline{a_{2}}\cdots\textup{\textcentoldstyle}_{1}\overline{a_{d}}\textup{\textcentoldstyle}_{1}$ where $\overline{0}=1$ and $\overline{1}=11$ (e.g., $\mathrm{VG}_{1}(011)=\textup{\textcentoldstyle}_{1}1\textup{\textcentoldstyle}_{1}11\textup{\textcentoldstyle}_{1}11\textup{\textcentoldstyle}_{1}$). Similarly, let $\mathrm{VG}_{2}(b)=\textup{\textcentoldstyle}_{2}\overline{b_{1}}\textup{\textcentoldstyle}_{2}\overline{b_{2}}\cdots\textup{\textcentoldstyle}_{2}\overline{b_{d}}\textup{\textcentoldstyle}_{2}$ for $b\in B$. Let
$w=\left(\bigodot_{a\in A}\textup{\#}\mathrm{VG}_{1}(a)\right)\textup{\textdollaroldstyle}\left(\bigodot_{b\in B}\mathrm{VG}_{2}(b)^{R}\textup{\#}\right).$
Both the length of $w$ and the time required for this construction are $O(nd)$.
We then define an RSQ $r$ over the alphabet $\Sigma=\{1,\textup{\#},\textup{\textdollaroldstyle},\textup{\textcentoldstyle}_{1},\textup{\textcentoldstyle}_{2}\}$. We write $r?$ for $r+\lambda$, where $\lambda$ denotes the empty string. Let
$r$ $=[1\textup{\#}\textup{\textcentoldstyle}_{1}\textup{\textcentoldstyle}_{2}]^{*}\textup{\#}r_{d}\textup{\#}[1\textup{\#}\textup{\textcentoldstyle}_{1}\textup{\textcentoldstyle}_{2}]^{*},$
$r_{k}$ $=\textup{\textcentoldstyle}_{1}{(1(r_{k-1}?)(1?))}\{2\}\textup{\textcentoldstyle}_{2}\text{ for }k\in\{1,\dots,d\},$
$r_{0}$ $=\textup{\textcentoldstyle}_{1}\Sigma^{*}\textup{\textcentoldstyle}_{2}.$
Both the length of $r$ and the time required for this construction are $O(d)$. To establish the correctness of the reduction, it suffices to show the following claim:
Claim.
Let $a\in A$, $b\in B$ and $s$ be a string over $\Sigma$ where no substring of the form $\textup{\textcentoldstyle}_{2}\Sigma^{*}\textup{\textcentoldstyle}_{1}$ occurs. Then, $a$ and $b$ are orthogonal if and only if $r_{d}$ matches $\mathrm{VG}_{1}(a)s\mathrm{VG}_{2}(b)^{R}$.
Proof.
We prove by induction on $d$. The case $d=0$ is trivial. Suppose that $d\geq 1$ and $a$ and $b$ are orthogonal. Let $e=\textup{\textcentoldstyle}_{1}1r_{d-1}1\textup{\textcentoldstyle}_{2}\,+\,\textup{\textcentoldstyle}_{1}1r_{d-1}11\textup{\textcentoldstyle}_{2}\,+\,\textup{\textcentoldstyle}_{1}11r_{d-1}1\textup{\textcentoldstyle}_{2}$. Observe that $r_{d}$ is equivalent to
$\textup{\textcentoldstyle}_{1}(1(1?))^{2}\textup{\textcentoldstyle}_{2}\>+\>\textup{\textcentoldstyle}_{1}1r_{d-1}(1?)1r_{d-1}(1?)\textup{\textcentoldstyle}_{2}\>+\>\textup{\textcentoldstyle}_{1}(1r_{d-1}111\>+\>111r_{d-1}(1?)\>+\>11r_{d-1})\textup{\textcentoldstyle}_{2}+e$ ($\ast$)
and therefore $r_{d}$ matches strings that $e$ matches. We write $u$ for $\mathrm{VG}_{1}(a)s\mathrm{VG}_{2}(b)^{R}$ and $u^{\prime}$ for $\textup{\textcentoldstyle}_{1}\overline{a_{2}}\cdots\textup{\textcentoldstyle}_{1}\overline{a_{d}}\,\textup{\textcentoldstyle}_{1}\,s\,\textup{\textcentoldstyle}_{2}\,\overline{b_{d}}^{R}\textup{\textcentoldstyle}_{2}\cdots\overline{b_{2}}^{R}\textup{\textcentoldstyle}_{2}$. By the induction hypothesis, $r_{d-1}$ matches $u^{\prime}$. From $a_{1}b_{1}=0$, it follows that $u$ is one of $\textup{\textcentoldstyle}_{1}1u^{\prime}1\textup{\textcentoldstyle}_{2}$, $\textup{\textcentoldstyle}_{1}1u^{\prime}11\textup{\textcentoldstyle}_{2}$ and $\textup{\textcentoldstyle}_{1}11u^{\prime}1\textup{\textcentoldstyle}_{2}$. Hence $e$ matches $u$ and so does $r_{d}$.
Conversely, suppose that $r_{d}$ matches $u$. We claim that so does $e$. This implies that $(\overline{a_{1}},\overline{b_{1}})\in\{(1,1),(1,11),(11,1)\}$ and hence $a_{1}b_{1}=0$. This completes the proof, because $a_{i}b_{i}=0$ for $i\in\{2,\cdots,d\}$ by the induction hypothesis and thus $a$ and $b$ are orthogonal. It suffices to show that no term in ($\ast$4.3) other than $e$ matches $u$. It is straightforward to see that neither $\textup{\textcentoldstyle}_{1}(1(1?))^{2}\textup{\textcentoldstyle}_{2}$ nor $\textup{\textcentoldstyle}_{1}(1r_{d-1}111+111r_{d-1}(1?)+11r_{d-1})\textup{\textcentoldstyle}_{2}$ matches $u$. Moreover, if $\textup{\textcentoldstyle}_{1}1r_{d-1}(1?)1r_{d-1}(1?)\textup{\textcentoldstyle}_{2}$ were to match $u$, then $u$ would have a substring of the form $\textup{\textcentoldstyle}_{2}(1?)1\textup{\textcentoldstyle}_{1}$ because $r_{d-1}$ only matches strings that begin with $\textup{\textcentoldstyle}_{1}$ and end with $\textup{\textcentoldstyle}_{2}$. This contradicts the fact that $\mathrm{VG}_{1}(a)$ contains no $\textup{\textcentoldstyle}_{2}$, $\mathrm{VG}_{2}(b)$ contains no $\textup{\textcentoldstyle}_{1}$, and no substring of the form $\textup{\textcentoldstyle}_{2}\Sigma^{*}\textup{\textcentoldstyle}_{1}$ occurs in $s$. ∎
Without loss of generality, we may assume that $2-\varepsilon\geq 1$. Using the RSQ matching algorithm in the hypothesis, we can solve the OV problem with $d=c\log{n}$ for any constant $c\geq 1$ in time
$O(nd+d+(nd)^{2-\varepsilon}\operatorname{poly}(d))=O(n^{2-\varepsilon}\operatorname{poly}(d))=O(n^{2-\varepsilon/2}).\qed$
Remark 9.
As in Remark 6, we can avoid using the Kleene star in the construction of $r$ by adopting a technique used in Lemma 2 of [47]. Let $e^{*}$ be a starred subexpression of $r$. Instead of $e^{*}$, we use ${(\cdots({({e?}\{2\})}\{2\})\cdots)}\{2\}$ where squares are nested $\lceil\log_{2}{|w|}\rceil$ times. Then, $r$ matches $w$ if and only if it matches $w$ after this change, because the matching is determined by which subexpressions of $r$ match which substrings of $w$ and the lengths of the latter are at most $|w|$. Although both the length of $r$ and the time required for the construction increase to $O(d+\log{n})$, the resulting lower bound remains unchanged because we allow $d$ to be logarithmic in $n$. Thus, Lemma 8 holds even when restricted to star-free RSQs.
5 Clique to ERE Matching
In this section, we prove Theorem 4. ERE [29] extends regular expressions with the complement construct $\lnot r$. The language of $\lnot r$ is the complement of the language of $r$, that is, the set of strings over the alphabet that are not matched by $r$. For example, $\lnot((aa)^{*})$ is equivalent to $a(aa)^{*}$ over the alphabet $\{a\}$. We use the intersection $r_{1}\cap r_{2}$ as a shorthand for $\lnot(\lnot r_{1}+\lnot r_{2})$.
The $k$-Clique problem asks whether a given undirected graph $G$ has a $k$-clique (i.e., the complete graph on $k$ vertices) as a subgraph. For a fixed $k$, a naive algorithm solves $k$-Clique in $O(n^{k})$ time. Abboud et al. [4] have proposed a polylogarithmically faster combinatorial algorithm for the problem.
Hypothesis (Combinatorial $k$-Clique Hypothesis, see [2, 4]).
For any integer $k\geq 3$ and any constant $\varepsilon>0$, no combinatorial algorithm can solve $k$-Clique on a graph of $n$ vertices in $O(n^{k-\varepsilon})$ time.
Using fast matrix multiplication, Nešetřil and Poljak showed a faster algorithm that runs in $O(n^{\omega\lceil k/3\rceil})$ time [42]. Here, $2\leq\omega<2.3714$ [7] denotes the exponent of square matrix multiplication.
Hypothesis ($k$-Clique Hypothesis, see [2, 59]).
For any integer $k\geq 3$ and any constant $\varepsilon>0$, no algorithm can solve $k$-Clique on a graph of $n$ vertices in $O(n^{\omega k/3-\varepsilon})$ time.
We show a combinatorial reduction from $3k$-Clique to ERE matching by adopting the approach of Abboud et al. [2] who showed a combinatorial reduction from $3k$-Clique to the recognition problem for context-free languages. Their reduction lists all $k$-cliques in a given graph and checks if there are three disjoint $k$-cliques that are fully connected with each other using a CFL recognition algorithm.
Lemma 10.
[…]
Proof.
We construct a reduction from $3k$-Clique to ERE matching as follows. Given a graph $G$ of $n$ vertices, we encode each vertex $v$ as a binary string $\overline{v}$ of fixed length $d=O(\log{n})$ and list all $k$-cliques $t_{1},\dots,t_{l}$ in $G$. We then encode them into a string $w$ as follows. For each $k$-clique $t=\{v_{1},\dots,v_{k}\}$, we define the following gadgets:
$\mathrm{VG}(t)$ $=\overline{v_{1}}\textup{\%}\overline{v_{2}}\textup{\%}\cdots\overline{v_{k}}\textup{\%},$
$\mathrm{LG}(t)$ $=\textup{\%}\mathrm{NG}(v_{1})\textup{\%}\mathrm{NG}(v_{2})\cdots\textup{\%}\mathrm{NG}(v_{k}),$
$\mathrm{NG}(v)$ $=\textup{\%}\overline{u_{1}}^{R}\textup{\%}\overline{u_{2}}^{R}\cdots\textup{\%}\overline{u_{p}}^{R}\quad\text{where $u_{1},\dots,u_{p}$ are the non-adjacent vertices of $v$}.$
For the sake of intuition, we name $\mathrm{VG}(t)$ a vertex gadget, $\mathrm{LG}(t)$ a list gadget, and $\mathrm{NG}(v)$ a non-adjacent vertex gadget. Remark that a list gadget $\mathrm{LG}(t)$ lists the non-adjacent vertices of each vertex of $t$. We then let
$w=\left(\bigodot_{i=1}^{l}\textup{\#}\mathrm{VG}(t_{i})\right)\textup{\#}\textup{\textdollaroldstyle}\left(\bigodot_{i=1}^{l}\textup{\#}\mathrm{LG}(t_{i})\textup{\textcentoldstyle}\mathrm{VG}(t_{i})\right)\textup{\#}\textup{\textdollaroldstyle}\left(\bigodot_{i=1}^{l}\textup{\#}\mathrm{LG}(t_{i})\right)\textup{\#}.$
Both the length of $w$ and the time required for the construction are $O(n^{k+1}\log{n})$.
[…]
$r$ $=[01\textup{\%}\textup{\#}]^{*}\textup{\#}((\lnot r_{\mathrm{disconn}}\textup{\textcentoldstyle}\lnot r_{\mathrm{disconn}})\cap\lnot r_{\mathrm{disconn}})\textup{\#}[01\textup{\%}\textup{\#}]^{*},$
$r_{\mathrm{disconn}}$ $=[01\textup{\%}]^{*}(r_{\mathrm{eq}}\cap[01]r_{\mathrm{eq}}[01]\cap\cdots\cap[01]^{d-1}r_{\mathrm{eq}}[01]^{d-1})[01\textup{\%}]^{*},$
$r_{\mathrm{eq}}$ $=0\Sigma^{*}0+1\Sigma^{*}1.$
Both the length of $r$ and the time required for this construction are $O(d^{2})=O(\log^{2}{n})$. To prove the correctness of the reduction, we need the following claim:
Claim.
Let $t$ and $t^{\prime}$ be two $k$-cliques in $G$ and $s\in\Sigma^{*}$. Then, $r_{\mathrm{disconn}}$ matches $\mathrm{VG}(t)\textup{\#}s\textup{\#}\mathrm{LG}(t^{\prime})$ if and only if $t\cup t^{\prime}$ is not a $2k$-clique in $G$.
Proof.
Suppose that $\mathrm{VG}(t)\textup{\#}s\textup{\#}\mathrm{LG}(t^{\prime})\in r_{\mathrm{disconn}}$. We can decompose it as $s_{1,0}s_{1,1}\textup{\#}s\textup{\#}s_{2,0}s_{2,1}$ where $\mathrm{VG}(t)=s_{1,0}s_{1,1},\mathrm{LG}(t^{\prime})=s_{2,0}s_{2,1}$ and $s_{1,1}\textup{\#}s\textup{\#}s_{2,0}\in r_{\mathrm{eq}}\cap[01]r_{\mathrm{eq}}[01]\cap\cdots\cap[01]^{d-1}r_{\mathrm{eq}}[01]^{d-1}$. From this and the definition of $r_{\mathrm{eq}}$, we observe that the middle substring $s_{1,1}\textup{\#}s\textup{\#}s_{2,0}$ has a binary encoding $\overline{v}$ of a vertex $v$ of $t$ as a prefix such that the reverse $\overline{v}^{R}$ is a member of $\mathrm{NG}(v^{\prime})$ for some vertex $v^{\prime}$ of $t^{\prime}$. This implies that $v$ is not adjacent to $v^{\prime}$ and $t\cup t^{\prime}$ does not form a $2k$-clique in $G$. The proof of the other direction is analogous and omitted. ∎
In what follows, we show that $G$ has a $3k$-clique if and only if $r$ matches $w$. For the only if direction, fix a $3k$-clique in $G$ and decompose it into three $k$-cliques $t_{1},t_{2}$ and $t_{3}$. We can decompose $w$ as
$w=w_{1,0}\textup{\#}\mathrm{VG}(t_{1})\textup{\#}w_{1,1}\textup{\textdollaroldstyle}w_{2,0}\textup{\#}\mathrm{LG}(t_{2})\textup{\textcentoldstyle}\mathrm{VG}(t_{2})\textup{\#}w_{2,1}\textup{\textdollaroldstyle}w_{3,0}\textup{\#}\mathrm{LG}(t_{3})\textup{\#}w_{3,1},$
where $w_{1,0},w_{1,1},w_{3,0},w_{3,1}\in[01\textup{\%}\textup{\#}]^{*}$ and $w_{2,0},w_{2,1}\in[01\textup{\%}\textup{\#}\textup{\textcentoldstyle}]^{*}$. Thus, it suffices to show that
  • $\mathrm{VG}(t_{1})\textup{\#}w_{1,1}\textup{\textdollaroldstyle}w_{2,0}\textup{\#}\mathrm{LG}(t_{2})\notin r_{\mathrm{disconn}}$,
  • $\mathrm{VG}(t_{2})\textup{\#}w_{2,1}\textup{\textdollaroldstyle}w_{3,0}\textup{\#}\mathrm{LG}(t_{3})\notin r_{\mathrm{disconn}}$ and
  • $\mathrm{VG}(t_{1})\textup{\#}w_{1,1}\textup{\textdollaroldstyle}w_{2,0}\textup{\#}\mathrm{LG}(t_{2})\textup{\textcentoldstyle}\mathrm{VG}(t_{2})\textup{\#}w_{2,1}\textup{\textdollaroldstyle}w_{3,0}\textup{\#}\mathrm{LG}(t_{3})\notin r_{\mathrm{disconn}}$,
and these follow from the previous claim because $t_{1}\cup t_{2}$, $t_{2}\cup t_{3}$ and $t_{1}\cup t_{3}$ are all $2k$-cliques in $G$.
For the if direction, suppose that $r$ matches $w$. The delimiters # and \$ enforce that the subexpression $(\lnot r_{\mathrm{disconn}}\textup{\textcentoldstyle}\lnot r_{\mathrm{disconn}})\cap\lnot r_{\mathrm{disconn}}$ of $r$ matches a substring $s$ of $w$ of the form $s=\mathrm{VG}(t_{1})\textup{\#}w_{1}\textup{\textdollaroldstyle}w_{2}\textup{\textdollaroldstyle}w_{3}\textup{\#}\mathrm{LG}(t_{3})$ for some $k$-cliques $t_{1}$ and $t_{3}$ in $G$, and some strings $w_{1},w_{3}\in[01\textup{\%}\textup{\#}]^{*}$ and $w_{2}\in[01\textup{\%}\textup{\#}\textup{\textcentoldstyle}]^{*}$. From $s\notin r_{\mathrm{disconn}}$ and the previous claim, $t_{1}\cup t_{3}$ forms a $2k$-clique in $G$. Moreover, from $s\in\lnot r_{\mathrm{disconn}}\textup{\textcentoldstyle}\lnot r_{\mathrm{disconn}}$ and the use of the delimiter ¢, we can write $w_{2}$ as $w_{2,0}\textup{\#}\mathrm{LG}(t_{2})\textup{\textcentoldstyle}\mathrm{VG}(t_{2})\textup{\#}w_{2,1}$ for some $k$-clique $t_{2}$ in $G$ and strings $w_{2,0},w_{2,1}\in[01\textup{\%}\textup{\#}\textup{\textcentoldstyle}]^{*}$, and
  • $\mathrm{VG}(t_{1})\textup{\#}w_{1}\textup{\textdollaroldstyle}w_{2,0}\textup{\#}\mathrm{LG}(t_{2})\notin r_{\mathrm{disconn}}$ and
  • $\mathrm{VG}(t_{2})\textup{\#}w_{2,1}\textup{\textdollaroldstyle}w_{3}\textup{\#}\mathrm{LG}(t_{3})\notin r_{\mathrm{disconn}}$.
Applying the previous claim, $t_{1}\cup t_{2}$ and $t_{2}\cup t_{3}$ form $2k$-cliques in $G$. Hence, $G$ has a $3k$-clique $t_{1}\cup t_{2}\cup t_{3}$ and we are done.
Therefore, using the ERE matching algorithm in the hypothesis, we can solve $3k$-Clique in time
$O(n^{k+1}\log{n}+\log^{2}{n}+T(n^{k+1}\log{n})\operatorname{poly}(\log^{2}{n}))=\tilde{O}(T(n^{k+1})).\qed$
Remark 11.
[…]
The following corollaries establish Theorem 4.
Corollary 12.
Let $\omega$ be the exponent of square matrix multiplication. If ERE matching on a string of length $n$ and an ERE of length $m$ can be solved in $O(n^{\omega-\varepsilon}\operatorname{poly}(m))$ time for some $\varepsilon>0$, then $k$-Clique on a graph of $n$ vertices can be solved in $O(n^{\omega k/3-\varepsilon^{\prime}})$ time for some sufficiently large integer $k$ and some $\varepsilon^{\prime}>0$, and the $k$-Clique Hypothesis is false.
Proof.
Without loss of generality, we may assume that $\omega-\varepsilon\geq 1$. Apply Lemma 10 with $T(n)=n^{\omega-\varepsilon}$ and sufficiently large $k$. Then, we can solve $3k$-Clique of $n$ vertices in time $\tilde{O}(n^{(k+1)(\omega-\varepsilon)})=O(n^{\omega k-\varepsilon^{\prime}})$ for some $\varepsilon^{\prime}>0$. ∎
Corollary 13.
If ERE matching on a string of length $n$ and an ERE of length $m$ can be solved combinatorially in $O(n^{3-\varepsilon}\operatorname{poly}(m))$ time for some $\varepsilon>0$, then $k$-Clique on a graph of $n$ vertices can be solved combinatorially in $O(n^{k-\varepsilon^{\prime}})$ time for some sufficiently large integer $k$ and some $\varepsilon^{\prime}>0$, and the Combinatorial $k$-Clique Hypothesis is false.
Proof.
The proof follows by replacing $\omega$ with $3$ in the proof of Corollary 12. ∎
Together with Theorem 4, the following lemma implies the optimality of the Hopcroft and Ullman’s algorithm in the sense described in Section 1.2. That is, the algorithm is combinatorially optimal and optimal with fast matrix multiplication in the sense that the exponents of $n$ in the respective running times cannot be shaved by any constant while preserving the running times to be constant-degree polynomial in $m$.
Lemma 14.
Let $T(n,m)$ be a function satisfying $\liminf_{n,m\to\infty}T(n,m)>0$ (e.g., $T(n,m)\geq n+m$) and $c>0$ be a constant. Then, $T(n,m)$ is $O(n^{c-\varepsilon}m^{l})$ for some constants $\varepsilon>0$ and $l\geq 0$ if and only if there exist constants $\varepsilon>0$ and $l\geq 0$ such that $T(n,m)$ is $O(n^{c-\varepsilon})$ for any fixed $m$ and $O(m^{l})$ for any fixed $n$.
Proof.
The only if direction is immediate. For the converse, let $\varepsilon$ and $l$ be as in the statement. By the first condition, there exists a constant $\delta>0$ such that $T(n,m)\geq\delta$ for all sufficiently large $n$ and $m$. Because $T(n,m)=O(n^{c-\varepsilon})$ for any fixed $m$, we have $T(n,m)\leq n^{c-\varepsilon/2}$ for all sufficiently large $n$. Similarly, because $T(n,m)=O(m^{l})$ for any fixed $n$, we have $T(n,m)\leq m^{l+1}$ for all sufficiently large $m$. Therefore,
$T(n,m)\leq T(n,m)^{2}/\delta\leq(1/\delta)n^{c-\varepsilon/2}m^{l+1}$
for all sufficiently large $n$ and $m$. ∎
We note that Lemma 14 also applies to Theorems 1, 2 and 3. Therefore, unless OVC fails, neither 1-rewb matching, RSQ matching nor semi-ERE matching can be solved within time both truly subquadratic in $n$ and constant-degree polynomial in $m$.
6 Conclusion
In this paper, we have studied the hardness of regular expression matching with four well-known extensions, backreference, squaring, intersection and complement. We have shown that (i) neither 1-rewb matching, RSQ matching nor semi-ERE matching can be solved in $O(n^{2-\varepsilon}\operatorname{poly}(m))$ time for any $\varepsilon>0$ under the Orthogonal Vectors Conjecture (Theorems 1, 2 and 3), and that (ii) ERE matching cannot be solved in $O(n^{\omega-\varepsilon}\operatorname{poly}(m))$ time for any $\varepsilon>0$ under the $k$-Clique Hypothesis and it cannot be solved combinatorially in $O(n^{3-\varepsilon}\operatorname{poly}(m))$ time for any $\varepsilon>0$ under the Combinatorial $k$-Clique Hypothesis (Theorem 4). As mentioned in the introduction, our results entail a number of interesting theoretical and practical consequences. For instance, our results provide evidence that unlike the case of lookaround, $O(nm)$-time matching algorithms for regular expressions extended with these features (in their full generality) are unlikely to exist, and that the ERE matching algorithm by Hopcroft and Ullman [29] was already optimal in the sense described in Section 1.2.
A possible direction for future work is to close the gap between the upper and lower bounds of 1-rewb matching, RSQ matching and semi-ERE matching. Regarding upper bounds for 1-rewb matching, as mentioned in Section 2, Nogami and Terauchi [45] recently proposed an $O(n^{2}m^{2})$-time algorithm for a certain subclass of 1-rewbs. If the general 1-rewb matching can also be solved in the same time complexity (or up to a polynomial factor in $m$), Theorem 1 provides evidence under OVC that there can be no further significant improvement in $n$ unless we allow a superpolynomial blowup in $m$. However, generalizing their algorithm to all 1-rewbs does not seem straightforward, as the algorithm makes non-trivial uses of the special structure of the 1-rewb subclass that it considers. Another possible direction for future work is to prove lower bounds on the matching problem with bounded repetition of constant nesting depth. This would provide further insights on the hardness of bounded repetition in practice, where it is typically not deeply nested.